# Umlaut information

### v1

@@ III.4 Example: Joint Gaussian distributions
$V=\begin{pmatrix}V_{XX}&V_{XY}\\ V_{XY}^{\intercal}&V_{YY}\end{pmatrix},$ (63)
@@
$P_{XY}(x,y)=\frac{e^{-\frac{1}{2}(r-m)^{\intercal}V^{-1}(r-m)}}{(2\pi)^{(n+k)/2}\sqrt{\det V}}.$ (64)
@@
$\ddot{P}_{Y}(y)=\frac{1}{(2\pi)^{k/2}\sqrt{\det(V/V_{XX})}}\,e^{-\frac{1}{2}(y-m_{Y})^{\intercal}(V/V_{XX})^{-1}(y-m_{Y})},$ (65)
@@
$U(X;Y)=\frac{1}{2}\log\frac{\det V}{\det\left(V_{XX}\oplus(V/V_{XX})\right)}+\frac{1}{2}\operatorname{Tr}\left[V^{-1}\left(V_{XX}\oplus(V/V_{XX})\right)\right]-\frac{n+k}{2}.$ (67)
@@
$V^{-1}=\begin{pmatrix}(V/V_{YY})^{-1}&Z\\ Z^{\intercal}&(V/V_{XX})^{-1}\end{pmatrix},$ (68)
@@
$P_{X}(x)=\frac{e^{-\frac{1}{2}(x-m_{X})^{\intercal}V_{XX}^{-1}(x-m_{X})}}{(2\pi)^{n/2}\sqrt{\det V_{XX}}}.$ (70)
@@
$\ddot{P}_{Y}(y)$ $\propto\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right)$ (71)
@@
$\sum_{x\in\mathcal{X}}\frac{1}{2}(r-m)^{\intercal}$ $V^{-1}(r-m)\frac{e^{-\frac{1}{2}(x-m_{X})^{\intercal}V_{XX}^{-1}(x-m_{X})}}{(2\pi)^{n/2}\sqrt{\det V_{XX}}}$ (72)
@@
$\ddot{P}_{Y}(y)=\frac{1}{(2\pi)^{k/2}\sqrt{\det(V/V_{XX})}}\,e^{-\frac{1}{2}(y-m_{Y})^{\intercal}(V/V_{XX})^{-1}(y-m_{Y})}.$ (73)
@@
$U(X;Y)=\frac{1}{2}\log\frac{\det V}{\det\left(V_{XX}\oplus(V/V_{XX})\right)}+\frac{1}{2}\operatorname{Tr}\left[V^{-1}\left(V_{XX}\oplus(V/V_{XX})\right)\right]-\frac{n+k}{2},$ (75)
@@ IV.1 Definition and basic properties
Let us introduce the notion of umlaut information of a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, i.e. a stochastic matrix $[\pazocal{W}(y|x)]_{x\in\mathcal{X},y\in\mathcal{Y}}$. Let us consider, as the input of the channel, a random variable $X$ taking values in $\mathcal{X}$ and with distribution $P_{X}$. The corresponding output will therefore be a random variable $Y$ taking values in $\mathcal{Y}$ with distribution
$P_{Y}(y)=\sum_{x\in\mathcal{X}}\pazocal{W}(y|x)P_{X}(x).$ (76)
@@
$P_{XY}(x,y)=\pazocal{W}(y|x)P_{X}(x)\,.$ (77)
Now we are interested in studying the relative entropy between the pair $(X,\tilde{Y})$ and the pair $(X,Y)$, where $\tilde{Y}$ is a random variable (independent of $X$) taking values in $\mathcal{Y}$ that minimises such divergence (see Figure 1). Then we maximise the result among the possible probability distributions of the input, as in the following definition.
@@
$U_{\alpha}(\pazocal{W})$ $\coloneqq\max_{P_{X}}U_{\alpha}(X;Y)=\max_{P_{X}}\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY}).$ (82)
As an immediate consequence of Lemma 5 we can see that the $\alpha$-Rényi umlaut information of a channel converges to the umlaut information.
@@
$\limsup_{\alpha\to 1^{-}}U_{\alpha}(\pazocal{W})=U(\pazocal{W}).$ (83)
IV.2 Zero-rate limit of the sphere-packing bound

### v2

@@ III.5 Example: Joint Gaussian distributions
$\hskip 0.0ptV=\begin{pmatrix}V_{XX}&V_{XY}\\ V_{XY}^{\intercal}&V_{YY}\end{pmatrix}$, (83)
@@
$\hskip 0.0ptP_{XY}(x,y)=\frac{e^{-\frac{1}{2}(r-m)^{\intercal}V^{-1}(r-m)}}{(2\pi)^{(n+k)/2}\sqrt{\det V}}.$ (84)
@@
$\hskip 0.0pt\ddot{P}_{Y}(y)=\frac{1}{(2\pi)^{k/2}\sqrt{\det(V/V_{XX})}}\,e^{-\frac{1}{2}(y-m_{Y})^{\intercal}(V/V_{XX})^{-1}(y-m_{Y})},$ (85)
@@
$\hskip 0.0ptU(X;Y)=\frac{1}{2}\log\frac{\det V}{\det\left(V_{XX}\oplus(V/V_{XX})\right)}+\frac{1}{2}\operatorname{Tr}\left[V^{-1}\left(V_{XX}\oplus(V/V_{XX})\right)\right]-\frac{n+k}{2}.$ (87)
@@
$\hskip 0.0ptV^{-1}=\begin{pmatrix}(V/V_{YY})^{-1}&Z\\ Z^{\intercal}&(V/V_{XX})^{-1}\end{pmatrix}$, (88)
@@
$\hskip 0.0ptP_{X}(x)=\frac{e^{-\frac{1}{2}(x-m_{X})^{\intercal}V_{XX}^{-1}(x-m_{X})}}{(2\pi)^{n/2}\sqrt{\det V_{XX}}}.$ (90)
@@
$\hskip 0.0pt\ddot{P}_{Y}(y)$ $\propto\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right)$ (91)
@@
$\hskip 0.0pt\sum_{x\in\mathcal{X}}\frac{1}{2}(r-m)^{\intercal}$ $V^{-1}(r-m)\frac{e^{-\frac{1}{2}(x-m_{X})^{\intercal}V_{XX}^{-1}(x-m_{X})}}{(2\pi)^{n/2}\sqrt{\det V_{XX}}}$ (92)
@@
$\hskip 0.0pt\ddot{P}_{Y}(y)=\frac{1}{(2\pi)^{k/2}\sqrt{\det(V/V_{XX})}}\,e^{-\frac{1}{2}(y-m_{Y})^{\intercal}(V/V_{XX})^{-1}(y-m_{Y})}.$ (93)
@@
$\hskip 0.0ptU(X;Y)=\frac{1}{2}\log\frac{\det V}{\det\left(V_{XX}\oplus(V/V_{XX})\right)}+\frac{1}{2}\operatorname{Tr}\left[V^{-1}\left(V_{XX}\oplus(V/V_{XX})\right)\right]-\frac{n+k}{2},$ (95)
@@ IV.1 Definition and basic properties
Let us introduce the notion of umlaut information of a discrete memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, i.e. a stochastic matrix $\pazocal{W}(y|x)]_{x\in\mathcal{X},y\in\mathcal{Y}}$. Let us consider, as the input of the channel, a random variable $X$ taking values in $\mathcal{X}$ and with distribution $P_{X}$. The corresponding output will therefore be a random variable $Y$ taking values in $\mathcal{Y}$ with distribution
$\hskip 0.0ptP_{Y}(y)=\sum_{x\in\mathcal{X}}\pazocal{W}(y|x)P_{X}(x).$ (96)
@@
$\hskip 0.0ptP_{XY}(x,y)=\pazocal{W}(y|x)P_{X}(x)\,.$ (97)
[…]
@@
IV.2 An explicit formula and additivity
Convention 17.
According to the usual conventions $0\log 0=0$ and $\exp(-\infty)=0$, we introduce the following notation. Let $\pazocal{W}$ be a discrete memoryless channel from $\mathcal{X}$ to $\mathcal{Y}$ and let $P_{X}\in\mathcal{P}(\mathcal{X})$. Then we define
$\hskip 0.0pt\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)\coloneqq\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)\log\pazocal{W}(y|x)\right),$ (102)
Proposition 18.
$\hskip 0.0ptU(\pazocal{W})=-\log\min_{P_{X}}\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)$ (103)
according to Convention 17, or, equivalently,
$\hskip 0.0pt\exp\left[-U(\pazocal{W})\right]$ $=\min_{P_{X}}\sum_{y\in\mathcal{Y}}\prod_{x\in\mathcal{X}}\pazocal{W}(y|x)^{P_{X}(x)},$ (104)
where we set $0^{0}=1$. It also holds that
$\hskip 0.0ptU(\pazocal{W})=\min_{Q_{Y}}\max_{x\in\mathcal{X}}D\left(Q_{Y}\middle\|\pazocal{W}(\,\cdot\,|x)\right)=\min_{Q_{Y}}\max_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}Q_{Y}(y)\log\frac{Q_{Y}(y)}{\pazocal{W}(y|x)}.$ (105)
Before proving the proposition, we recall a minimax lemma.
Lemma 19 ([24, Theorem 5.2]).
Let $A$ be a compact, convex subset of a Hausdorff topological vector space $U$, and let $B$ be a convex subset of the linear space $V$. Let $f:A\times B\to(-\infty,+\infty]$ be lower semicontinuous on $A$ for all fixed $y\in B$, concave in the first variable, and convex in the second. Then
$\hskip 0.0pt\sup_{x\in A}\inf_{y\in B}f(x,y)=\inf_{y\in B}\sup_{x\in A}f(x,y).$ (106)
Proof of Proposition 18.
$\hskip 0.0ptU(\pazocal{W})$ $=\max_{P_{X}}\left(-H(P_{X})-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\left(P_{X}(x)\pazocal{W}(y|x)\right)\right)\right)$ (107)
$\hskip 0.0ptU(\pazocal{W})$ $=\max_{P_{X}}\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})$ (108)
The expression to be optimised is linear in $P_{X}$ and convex in $Q_{Y}$, since it is the sum of two convex functions. Furthermore, $Q_{Y}\mapsto D(P_{X}Q_{Y}\|P_{XY})$ is lower semicontinuous [25, Theorem 15]. We can therefore apply the minimax result in Lemma 19 to get
$\hskip 0.0ptU(\pazocal{W})$ $=\min_{Q_{Y}}\max_{P_{X}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}(y)\log\frac{Q_{Y}(y)}{\pazocal{W}(y|x)}$ (109)
which concludes the proof. ∎
Corollary 20 ((Additivity of the channel umlaut information)).
$\hskip 0.0pt\big{(}\pazocal{W}_{1}\times\pazocal{W}_{2})(y_{1},y_{2}|x_{1},x_{2})\coloneqq\pazocal{W}_{1}(y_{1}|x_{1})\pazocal{W}_{2}(y_{2}|x_{2})$ (110)
$\hskip 0.0ptU(\pazocal{W}_{1}\times\pazocal{W}_{2})=U(\pazocal{W}_{1})+U(\pazocal{W}_{2}).$ (111)
Proof.
By (103), we immediately see that
$\hskip 0.0ptU(\pazocal{W}_{1}\times\pazocal{W}_{2})$ $=-\min_{P_{X_{1}X_{2}}}\log\sum_{\begin{subarray}{c}y_{1}\in\mathcal{Y}_{1}\\ y_{2}\in\mathcal{Y}_{2}\end{subarray}}\exp\left(\sum_{\begin{subarray}{c}x_{1}\in\mathcal{X}_{1}\\ x_{2}\in\mathcal{X}_{2}\end{subarray}}P_{X_{1}X_{2}}(x_{1},x_{2})\left(\log\pazocal{W}_{1}(y_{1}|x_{1})+\log\pazocal{W}_{2}(y_{2}|x_{2})\right)\right)$ (112)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}-\min_{P_{X_{1}}}\log\sum_{y_{1}\in\mathcal{Y}_{1}}\exp\left(\sum_{x_{1}\in\mathcal{X}_{1}}P_{X_{1}}(x_{1})\log\pazocal{W}_{1}(y_{1}|x_{1})\right)$
$=U(\pazocal{W}_{1})+U(\pazocal{W}_{2})$
IV.3 Rényi $\alpha$-umlaut information of a channel
@@
$\hskip 0.0ptU_{\alpha}(\pazocal{W})$ $\coloneqq\max_{P_{X}}U_{\alpha}(X;Y)=\max_{P_{X}}\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY}).$ (113)
Proposition 22.
The minimax variational form for the channel umlaut information in (100) can be generalised to the family of Rényi $\alpha$-umlaut information of channels for $\alpha\in(0,1)\cup(1,\infty)$. More precisely, let $\pazocal{W}$ be a discrete memoryless channel from $\mathcal{X}$ to $\mathcal{Y}$. Then
$U_{\alpha}(\pazocal{W})=\min_{Q_{Y}}\max_{x\in\mathcal{X}}D_{\alpha}(Q_{Y}\|\pazocal{W}(\,\cdot\,|x))\,.$ (114)
Proof.
We start by explicitly writing
$\hskip 0.0ptU_{\alpha}(\pazocal{W})$ $=\max_{P_{X}}\min_{Q_{Y}}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)P_{X}^{1-\alpha}(x)$ (115)
$=\max_{P_{X}}\min_{Q_{Y}}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x).$
Let us first consider the case $\alpha\in(0,1)$. By monotonicity of the logarithm, we have
$\hskip 0.0ptU_{\alpha}(\pazocal{W})$ $=\frac{1}{\alpha-1}\log\min_{P_{X}}\max_{Q_{Y}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x).$ (116)
By concavity of $x\mapsto x^{\alpha}$, the map $Q_{Y}\mapsto\sum_{x,y}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}$ is concave, and by linearity in $P_{X}$ we can apply Sion’s minimax theorem to switch maximum and minimum:
$\hskip 0.0ptU_{\alpha}(\pazocal{W})$ $=\frac{1}{\alpha-1}\log\max_{Q_{Y}}\min_{P_{X}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)$ (117)
$=\frac{1}{\alpha-1}\log\max_{Q_{Y}}\min_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)$
$=\min_{Q_{Y}}\max_{x\in\mathcal{X}}\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}}Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x),$
and this proves the claim for $\alpha\in(0,1)$. Analogously, for $\alpha>1$ we have
$\hskip 0.0ptU_{\alpha}(\pazocal{W})$ $=\frac{1}{\alpha-1}\log\max_{P_{X}}\min_{Q_{Y}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)$ (118)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\frac{1}{\alpha-1}\log\min_{Q_{Y}}\max_{P_{X}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)$
$=\frac{1}{\alpha-1}\log\min_{Q_{Y}}\max_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x)$
$=\min_{Q_{Y}}\max_{x\in\mathcal{X}}\frac{1}{\alpha-1}\log\sum_{y\in\mathcal{Y}}Q_{Y}^{\alpha}(y)\pazocal{W}^{1-\alpha}(y|x),$
where in (i) we noticed that $x\mapsto x^{\alpha}$ is convex. This concludes the proof. ∎
Corollary 23.
For all $\alpha\in(0,1)\cup(1,\infty)$, the Rényi $\alpha$-umlaut information is additive under the tensor product of channels, namely
$\hskip 0.0ptU_{\alpha}(\pazocal{W}_{1}\times\pazocal{W}_{2})=U_{\alpha}(\pazocal{W}_{1})+U_{\alpha}(\pazocal{W}_{2})$ (119)
Proof.
Using the variational form established in Proposition 22, we have
$\hskip 0.0ptU_{\alpha}(\pazocal{W}_{1}\times\pazocal{W}_{2})$ $=\min_{Q_{Y_{1}Y_{2}}}\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}_{1}\\ x_{2}\in\mathcal{X}_{2}\end{subarray}}D_{\alpha}(Q_{Y_{1}Y_{2}}\|\pazocal{W}_{1}(\cdot|x_{1})\times\pazocal{W}_{2}(\cdot|x_{2}))$ (120)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{\leq}}\min_{Q_{Y_{1}}\times Q_{Y_{2}}}\max_{\begin{subarray}{c}x_{1}\in\mathcal{X}_{1}\\ x_{2}\in\mathcal{X}_{2}\end{subarray}}D_{\alpha}(Q_{Y_{1}}\times Q_{Y_{2}}\|\pazocal{W}_{1}(\cdot|x_{1})\times\pazocal{W}_{2}(\cdot|x_{2}))$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}\min_{Q_{Y_{1}}}\max_{x_{1}\in\mathcal{X}_{1}}D_{\alpha}(Q_{Y_{1}}\|\pazocal{W}_{1}(\cdot|x_{1}))+\min_{Q_{Y_{2}}}\max_{x_{2}\in\mathcal{X}_{2}}D_{\alpha}(Q_{Y_{2}}\|\pazocal{W}_{2}(\cdot|x_{2}))$
$=U_{\alpha}(\pazocal{W}_{1})+U_{\alpha}(\pazocal{W}_{2})\,$,
where in (i) we have chosen the ansatz $Q_{Y_{1}Y_{2}}=Q_{Y_{1}}Q_{Y_{2}}$ and in (ii) we have leveraged the additivity of the Rényi $\alpha$-relative entropy. Conversely,
$\hskip 0.0ptU_{\alpha}(\pazocal{W}_{1}\times\pazocal{W}_{2})$ $=\max_{P_{X_{1}X_{2}}}U_{\alpha}(X_{1},X_{2}\,;Y_{1},Y_{2})_{(\pazocal{W}_{1})_{Y_{1}|X_{1}}(\pazocal{W}_{2})_{Y_{2}|X_{2}}P_{X_{1}X_{2}}}$ (121)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{\geq}}\max_{P_{X_{1}}\times P_{X_{2}}}U_{\alpha}(X_{1},X_{2}\,;Y_{1},Y_{2})_{(\pazocal{W}_{1})_{Y_{1}|X_{1}}P_{X_{1}}\times(\pazocal{W}_{2})_{Y_{2}|X_{2}}P_{X_{2}}}$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(vi)}}}}{{=}}\max_{P_{X_{1}}\times P_{X_{2}}}\left(U_{\alpha}(X_{1}\,;Y_{1})_{(\pazocal{W}_{1})_{Y_{1}|X_{1}}P_{X_{1}}}+U_{\alpha}(X_{2}\,;Y_{2})_{(\pazocal{W}_{2})_{Y_{2}|X_{2}}P_{X_{2}}}\right)$
$=U_{\alpha}(\pazocal{W}_{1})+U_{\alpha}(\pazocal{W}_{2}),$
where in (iii) we have chosen the ansatz $P_{X_{1}X_{2}}=P_{X_{1}}P_{X_{2}}$ and in (vi) we have used the additivity of the Rényi $\alpha$-umlaut information for probability distributions (Proposition 10). ∎
Finally, as an immediate consequence of Lemma 9 we can see that the $\alpha$-Rényi umlaut information of a channel converges to the umlaut information.
@@
$\hskip 0.0pt\lim_{\alpha\to 1}U_{\alpha}(\pazocal{W})=U(\pazocal{W})\,.$ (122)
Proof.
Using Proposition 22, we have
$\hskip 0.0pt\lim_{\alpha\to 1^{+}}U_{\alpha}(\pazocal{W})=\min_{Q_{Y}}\inf_{\alpha>1}\max_{x\in\mathcal{X}}D_{\alpha}(Q_{Y}\|\pazocal{W}(\,\cdot\,|x))\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\min_{Q_{Y}}\max_{x\in\mathcal{X}}\inf_{\alpha>1}D_{\alpha}(Q_{Y}\|\pazocal{W}(\,\cdot\,|x))\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}U(\pazocal{W}),$ (123)
where in (i) we noticed that $\mathcal{X}$ is finite and that, by the monotonicity in $\alpha$ of the Rényi divergences and by Lemma 9, we have
$\hskip 0.0pt\inf_{\alpha>1}D_{\alpha}(Q_{Y}\|\pazocal{W}(\,\cdot\,|x))=\lim_{\alpha\to 1^{+}}D_{\alpha}(Q_{Y}\|\pazocal{W}(\,\cdot\,|x))=D(Q_{Y}\|\pazocal{W}(\,\cdot\,|x)),$ (124)
and in (ii) we used (100). Analogously, using Lemma 9 we obtain the complementary claim, i.e.
$\hskip 0.0pt\lim_{\alpha\to 1^{-}}U_{\alpha}(\pazocal{W})=\sup_{\alpha<1}\max_{P_{X}}U_{\alpha}(X;Y)_{\pazocal{W}_{Y|X}P_{X}}=\max_{P_{X}}\sup_{\alpha<1}U_{\alpha}(X;Y)_{\pazocal{W}_{Y|X}P_{X}}=U(\pazocal{W}).$ (125)
IV.4 Zero-rate limit of the sphere-packing bound